OpenStudy (patriciab):

simplify 94-24-2-2/5*1/2

OpenStudy (anonymous):

the answer is B

OpenStudy (patriciab):

do i start by following the order of operation

OpenStudy (anonymous):

Hard to tell what the order is without parentheses..

OpenStudy (anonymous):

Here, though, there doesn't seem to be much ambiguity. You can think of it as \[94-24-2-2÷5 \times 1÷2 \space \space or \space \space 94-24-2-\frac{2}{5} \times \frac{1}{2}\] and get the same answer.